OpenStudy (anonymous):

evaluate , please show work . y=2x^2+r if y=11r and r=5 , find x

OpenStudy (anonymous):

11r = 2x^2 + 5 55 = 2x^2 + 5 2x^2 = 50 x^2 = 25 x = -5, 5

OpenStudy (anonymous):

thanks a bunch it seems that i missed the substitution for the 11r !